OpenStudy (anonymous):

What is the equation of the line that passes through the point (4, 1) and is perpendicular to the line y = –2x + 4?

OpenStudy (anonymous):

A line perpendicular to this line y=-2x+4 has a slope of 1/2. Use y=mx+b and solve for b. 1=(1/2)*(4)+b 1=2+b b=1-2 = -1 so the equation of the line is y=1/2x-1